There exists a uniform electric field E=4×10^5V/m directed along negative x axis such that the electric potential at origin is 0.A charge of -200micro coulomb is placed at origin and a charge of +200micro coulomb is placed at (3m,0).The electrostatic potential energy of system is ....

Grade:12

1 Answers

Arun
4 years ago

Dipole moment of the two charges is

 

p = q × l

p = 200×10^-4 × 3

p = 1.2×10^-3

 

P.E. = pEcos∅

P.E. = 1.2×10^-3 × 4×10^5 [ cos∅ = 0°]

P.E. = 4.8×10^2 J